Determine whether 4s2 + 48t² + 6r - 9 is an odd integer for all r, s and t in Z.
Answer #1

An odd integer is the form of 2m +1, where mez 45 +48+² +66-9 = 45²+48 +60-10 + 1 = 2(25++ 24ť +3r - 5) +1 - 2 m +1 where m=